# Relationship between the ray variability and the pc-scale jet in the blazar 3C 454.3

### v1

@@ 3.1 Jet Structure and Kinematics
In our analysis of the VLBA maps at 43 GHz we characterized the core (Q0), assumed to be a stationary element situated at the eastern end of the jet, along with several other notable features (see Fig. 3). We observed a quasi-stationary component oscillating between 0.5 and 0.6 mas, which was visible during two different time intervals: Q8 between 2011 and 2014, and Q21 from 2016 to 2020. This component was previously observed and labeled St (e.g. Jorstad et al., 2005, 2013) or region C (e.g. Jorstad et al., 2017; Traianou et al., 2024). Additionally, we identified 20 moving components labeled Q1, Q2, and so on, up to Q24 (excluding Q8 and Q21, as already mentioned). Most of the modeled components are consistent with knots discussed in earlier studies. Fig. 3 presents four different epochs where we identified the component, Q3, which was ejected from the core in early 2010, corresponding with the time of the ray flare observed between September 2009 and January 2010. This component was previously identified as K09, as reported by Jorstad et al. (2013), and its passage through the core was associated with a multi-wavelength outburst in 3C 454.3 (León-Tavares et al., 2013; Jorstad et al., 2013; Amaya-AlmazÔn et al., 2021).
@@
Similarly, Q10, Q17, Q19, and Q18 seemed to align with the knots referenced by Traianou etĀ al. (2024) as the disappearing knots observed at 43 and 86 GHz. These four components ejected from the core in February 2014, July 2014, February 2015, and May 2015, respectively, seemed to vanish once they approached the quasi-stationary component Q21 (or region C in their study), as depicted in Fig. 4. They proposed that this disappearance might be attributed to the bending of the jet at that specific location, causing the plasma to shift toward the observer line of sight.
At this particular frequency, the emission is mainly dominated by Q0, with only occasional periods when this is not the case. For instance: in August 2010 Q3 reached a maximum flux of 16.9 Jy/beam (see Fig. 3), which corresponds to just over 170 of the core flux at the same time. Between 2014 and 2016, the flux was primarily dominated by the mobile component Q10, which surpassed the core flux by a factor of 7. In 2016, the emission was primarily dominated by the quasi-stationary component Q21, with an average flux twice that of the core for that year.
Fig. 2 presents the results of model fitting applied to 3C 454.3 at 15 GHz. The image reveals a complex jet structure oriented towards the northwest, exhibiting subtle variations over time. This finding is further supported by observations from August 2008 to the end of 2020, which consistently depict the source as core-dominated with a persistent one-sided jet.
The brightest component, designated K0, is followed by K1, a quasi-stationary jet feature located on average 0.4 mas away. Notably, K1 remained attached to the core from mid-2008 to 2013. By the end of 2020, a new component, K6, emerged. K6 originated from K1 initial position and moved away from K0, potentially linked to the ray emission, as discussed later. Given that we excluded the 15 GHz VLBA 2019 data from our analysis (due to the reported flux discrepancies), we cannot definitively confirm if K6 originated from the same K1 component observed earlier. In addition to the moving components and the core, we identified two other quasi-stationary features, K4 and K5, located at approximately 6.5 mas and 8.5 mas from the core, respectively.
@@
where is the proper motion in rad s-1, is the luminosity distance in m, is the source redshift, and is the speed of light in m s-1.
@@ 3.2 Correlation analysis
The connection between the ray state of blazars and their parsec-scale jet behavior is well established. Studies by León-Tavares etĀ al., 2011 suggest that the brightest ray flares in blazars might be triggered by the initial stages of millimeter-wavelength outbursts. To quantify the potential link between the variability of the parsec-scale jet knots in 3C 454.3 and its ray emission, we performed Spearmanās rank correlation tests. This non-parametric test assesses the strength and direction of a monotonic relationship between two variables, independent of their underlying distributions. To account for the possibility of rapid ray variations reported in previous studies (e.g., Ackermann etĀ al. 2016; Pittori etĀ al. 2018), we implemented an adaptive time binning approach. We calculated the ray fluxes within time bins centered on the exact observation times of the VLBA data. The bin sizes were iteratively increased, starting small and always maintaining the VLBA observation date as the center. This process continued until the test statistic (TS) for each bin exceeded a threshold of 25. We consider a given correlation as significant if the probability of obtaining a specific correlation coefficient due to statistical fluctuations is less than 5, that is, if p-value . This approach ensures we capture potential correlations while maintaining the temporal resolution necessary to detect rapid variability.
@@
To further test the robustness of the observed correlations and minimize potential biases from individual data points or small subsets, we performed a Jackknife analysis on each component. This analysis, which involves iteratively removing one data point and recalculating the correlation, demonstrated consistent results. The mean correlation coefficients and p-values obtained from the Jackknife analysis were in good agreement with those presented in Tables 8 and 9 in all cases. Also to assess the reliability of the presented correlations, we adopted the methodology described in Alexander (1997). We conducted 10,000 Monte Carlo simulations, generating simulated datasets with fluxes drawn randomly from within their respective error ranges. Spearman rank correlations were then computed for each simulated dataset. Subsequently, we performed Fisherās z-transform on the resulting correlation coefficients, calculating the median and standard deviation in the z-space. The median was then back-transformed to obtain a correlation coefficient, and asymmetric errors were estimated. The last column of Tables 8 and 9 presents the correlation coefficients and associated errors determined through this Monte Carlo approach, which are consistent with the Spearman correlations obtained from the original data.
@@
The results of our analysis are presented in Table 8, they reveal significant correlations between the ray emission and the 43 GHz flux densities of specific jet components found in the BEAM-ME maps. The core (Q0) shows the strongest correlation, with a coefficient () of 0.52. This suggests a possible link between core variability at 43 GHz and roughly 27 of the observed ray variability, as indicated by a Spearman rank correlation coefficient squared () of 0.27. Similarly, correlations with components Q3 and Q21 suggest that 26 and 18 of the ray variability, respectively, could be related to the 43 GHz emission in these components during their detection periods. Fig. 6 illustrates these relationships between the simultaneous ray fluxes and the 43 GHz flux densities of the core and components Q3 and Q21. The data points associated with Q8 and Q21 correspond to two distinct time periods when the quasi-stationary component was observed. This component has been extensively studied in the literature prior to this work (e.g. Pauliny-Toth etĀ al. 1987, Jorstad etĀ al. 2001) and before the Fermi era. Intriguingly, during the first epoch, Q8 exhibits no correlation with the gamma-ray variability, while Q21 shows a clear correlation. As illustrated in Fig. 4, numerous moving features intersect with Q21, whereas none pass through Q8. This observation aligns with models proposed in previous studies that attribute ray flares to shock-shock interactions between moving and stationary features (e.g. Marscher & Gear 1985; Fromm etĀ al. 2011). These findings suggest that the quasi-stationary component may act as a ray emission region, with its activity modulated by the occurrence of such interactions.
@@
Fig. 6, right panels, also depict the relationships between the ray fluxes and the 15 GHz flux densities of the core (K0) and components K6 and K8 from the MOJAVE maps. Table 9 summarizes the Spearman rank correlation results. The core, once again, exhibits a strong and significant correlation with ray emission, with a coefficient of 0.62. This implies that roughly 39 of the ray variations are likely linked to the core 15 GHz activity. Interestingly, a new jet component, K6, emerged near the end of the study period (early 2020). This component shows a significant anti-correlation with the ray flux, with a coefficient of , suggesting that nearly 40 of the ray variability during this specific time might be related to the ejection and 15 GHz emission of K6, even in the absence of a simultaneous flare. As suggested by PatiƱo-Ćlvarez etĀ al. (2018), one possible explanation for the observed behavior is that the jet has become less transparent to rays. This change in transparency might be caused by gamma-ray absorption due to triple pair production. The cross-section for both the inverse Compton scattering and the triple pair production are dependent on the energy of the electrons in the jet, when the Lorentz factors of individual electrons increase up to a certain threshold, the triple pair production becomes more prominent than the inverse Compton. This process would absorb rays, causing a flux decrease, while at the same time, increasing the synchrotron emission, due to the increase in the Lorentz factor, and the production of more leptons within the jet. However, K6 proximity to the core as seen in Fig. 2, near the spatial resolution limit, raises the possibility of flux contamination. Additionally, a quasi-stationary component, K8, was detected at approximately 53 pc from the core. Although correlation analysis was feasible with 11 epochs, more observations are essential to determine this relationship.
@@
This study offers novel insights into the ray emission in blazars. Our findings reveal the presence of multiple distinct regions within the jet, each contributing to the high-energy emission. Notably, some of these regions exhibit non-stationary behavior, displaying temporal variability. Furthermore, we have successfully located these emission regions within the jet structure. Previous multi-frequency studies of 3C 454.3 have revealed strong correlations between ray and optical emissions (Ogle etĀ al. 2011; León-Tavares etĀ al. 2013; VillicaƱa-Pedraza etĀ al. 2016; Nalewajko etĀ al. 2019; Amaya-AlmazĆ”n etĀ al. 2021; Fan etĀ al. 2021). These correlations strongly suggest that rays in blazars originate from leptonic processes rather than hadronic ones, as no known hadronic mechanism can simultaneously generate rays and significantly impact optical synchrotron emission. To distinguish between SSC and EC as the dominant leptonic process, the location of the emitting region within the jet is crucial. For regions located several parsecs away from the central engine, the density of external seed photons is typically lower than the synchrotron photon density, favoring SSC as the primary emission mechanism. This is supported by findings in 3C 279, where PatiƱo-Ćlvarez etĀ al. (2019) concluded that approximately 33% of the observed ray emission originates from SSC processes within a particular jet component, during the observable period of said component. In the core region of the blazar, while we can confirm a leptonic origin for the gamma-ray emission, differentiating between SSC and EC remains challenging. However, it is generally accepted that sufficiently energetic electrons exist in the core to power the observed ray emission. For non-core components (i.e., those located farther from the central pc), the correlation observed in simultaneous data suggests that the ray emission mechanism is SSC. The electron energy distribution within a knot is such that while lower-energy electrons dominate the radio band, higher-energy electrons capable of producing gamma-rays can co-exist.
These findings could guide the development of more comprehensive theoretical models that might account for the observed variability and spatial distribution of high-energy emissions.
@@ 4 Summary and conclusions
In this study, we analyzed the jet behavior and structure of the blazar 3C 454.3 using radio data from the VLBA-BU-BLAZAR, BEAM-ME, and MOJAVE monitoring programs (at frequencies of 43 GHz and 15 GHz). We investigated potential ray emission regions within the jet components identified in the VLBA maps. This included measuring the distances between these regions and the radio core and exploring possible emission mechanisms. Additionally, we examined whether these regions account for all the ray emission, or if, like in the case of 3C 279 (PatiƱo-Ćlvarez etĀ al., 2019), multiple regions are needed to describe it.
Our key findings include:
@@
Both 43 GHz and 15 GHz core emissions strongly correlate with ray emission. We observed a robust correlation between the 43 GHz core emission and ray emission, with a correlation coefficient of 0.52. This indicates that approximately 27 of the variability in ray emission is related to the variability in the core radio emission. Similarly, the flux of the 15 GHz core exhibits a strong and statistically significant correlation with the ray emission, contributing approximately 40 to the high-energy emission.
@@
The quasi-stationary component Q21 in the 43 GHz maps, visible between 2016 and 2021, at a projected distance of 4.6 pc (0.6 mas), strongly correlates with the ray flux, implying that it may contribute up to 18% of the high-energy emission. However, the quasi-stationary component observed during a different time interval, Q8, does not correlate. This variability in the correlation suggests that quasi-stationary components may exhibit intermittent gamma-ray emission.
@@
At 43 GHz, we detected a mobile component whose flux exhibits a robust correlation with the ray flux, contributing to 26 of the high-energy emission between 2010.18 and 2011.16. This component, denoted Q3, demonstrates an apparent velocity of c, observed at a projected distance ranging from 0.8 to 2.3 pc (0.1 to 0.3 mas).
@@
For ray emission regions located farther than one parsec from the central black hole (e.g., components Q21 and Q3), our findings strongly suggest synchrotron self-Compton (SSC) as the primary mechanism responsible for the high-energy radiation. This conclusion is supported by the observed correlation between the radio variability in these regions and the ray emission, indicating a nearly simultaneous variability.
@@
In the 15 GHz maps, a new jet component, K6, appeared in early 2020. It strongly anti-correlates with ray flux (correlation coefficient -0.65), suggesting a potential link to around 40 of the ray variability. However, K6 proximity to the core might distort its measured flux. Another component, K8, shows a similar anti-correlation with rays, but we believe more data is needed to confirm this behavior.
@@
The observed anti-correlations in the two 15 GHz components may be attributed to an increased Lorentz factor in the electrons of the ray emission region, which can enhance electron-positron triple pair-production, leading to greater ray absorption and a subsequent decrease in observed gamma-ray flux, as suggested by PatiƱo-Ćlvarez etĀ al. (2018). This would also increase the synchrotron emission, because the energy of the leptons is increasing, as well as the number of said leptons, thus creating the observed anti-correlation.
These findings are valuable on two fronts. First, they provide crucial data for refining theoretical models of blazar jets. Second, they open new avenues for investigation by pinpointing specific regions within the jet for further study. This comprehensive understanding might lead to a deeper grasp of the complex processes at play in blazars.

### v2

@@ 3.1 Jet structure and kinematics
In our analysis of the VLBA maps at 43 GHz, we characterized the core (Q0), which we assumed to be a stationary element situated at the eastern end of the jet, along with several other notable features (see Fig. 3). We observed a quasi-stationary component oscillating between 0.5 and 0.6 mas, which was visible during two different time intervals: Q8 between 2011 and 2014 and Q21 from 2016 to 2020. This component was previously observed and labeled as St (e.g., Jorstad et al., 2005, 2013) or region C (e.g., Jorstad et al., 2017; Traianou et al., 2024). Additionally, we identified 20 moving components, labeled as Q1, Q2, and so on up to Q24 (excluding Q8 and Q21, as already mentioned). Most of the modeled components are consistent with knots discussed in earlier studies. Figure 3 presents four different epochs where we identified component Q3, ejected from the core in early 2010. This ejection coincides with the ray flare observed between September 2009 and January 2010. Previously, Jorstad et al. (2013) identified this component as K09, and its passage through the core was associated with a multi-wavelength outburst in 3C 454.3 (León-Tavares et al., 2013; Jorstad et al., 2013; Amaya-AlmazÔn et al., 2021).
@@
Similarly, Q10, Q17, Q19, and Q18 seem to align with the knots referenced by Traianou etĀ al. (2024) as the disappearing knots observed at 43 and 86 GHz. These four components ejected from the core in February 2014, July 2014, February 2015, and May 2015, respectively, seemed to vanish once they approached the quasi-stationary component Q21 (or region C in their study), as depicted in Fig. 4. Traianou etĀ al. (2024) proposed that this disappearance might be attributed to the bending of the jet at that specific location, thus causing the plasma to shift toward the observerās line of sight.
At this particular frequency, the emission is mainly dominated by Q0, with only occasional periods when this is not the case. For instance, in August 2010, Q3 reached a maximum flux of 16.9 Jy/beam (see Fig. 3), which corresponds to just over 170 of the core flux at the same time. Between 2014 and 2016, the flux was primarily dominated by the mobile component Q10, which surpassed the core flux by a factor of seven. In 2016, the emission was primarily dominated by the quasi-stationary component Q21 and had an average flux twice that of the core for that year.
Figure 2 presents the results of model fitting applied to 3C 454.3 at 15 GHz. The image reveals a complex jet structure oriented toward the northwest that exhibits subtle variations over time. This finding is further supported by observations from August 2008 to the end of 2020, which consistently depict the source as core dominated with a persistent one-sided jet.
The brightest component, designated as K0, is followed by K1, a quasi-stationary jet feature located on average 0.4 mas away. Notably, K1 remained attached to the core from mid-2008 to 2013. By the end of 2020, a new component, K6, emerged. K6 originated from the initial position of K1 and moved away from K0, and it is potentially linked to the -ray emission, as we discuss later. Given that we excluded the 15 GHz VLBA 2019 data from our analysis (due to the reported flux discrepancies), we cannot definitively confirm if K6 originated from the same K1 component observed earlier. In addition to the moving components and the core, we identified two other quasi-stationary features, K4 and K5, located at approximately 6.5 mas and 8.5 mas from the core, respectively.
@@
where is the proper motion in radians per second, is the luminosity distance in meters, is the source redshift, and is the speed of light in meters per second.
@@ 3.2 Correlation analysis
The connection between the -ray state of blazars and their parsec-scale jet behavior is well established. Studies by León-Tavares etĀ al., 2011 suggested that the brightest -ray flares in blazars might be triggered by the initial stages of millimeter-wavelength outbursts. To quantify the potential link between the variability of the parsec-scale jet knots in 3C 454.3 and its -ray emission, we performed Spearmanās rank correlation tests. This non-parametric test assesses the strength and direction of a monotonic relationship between two variables, independent of their underlying distributions. To account for the possibility of rapid -ray variations reported in previous studies (e.g., Ackermann etĀ al. 2016; Pittori etĀ al. 2018), we implemented an adaptive time binning approach. We calculated the -ray fluxes within time bins centered on the exact observation times of the VLBA data. The bin sizes were iteratively increased, starting small and always maintaining the VLBA observation date as the center. This process continued until the test statistic for each bin exceeded a threshold of 25. We consider a given correlation as significant if the probability of obtaining a specific correlation coefficient due to statistical fluctuations is less than 5 (i.e., p-value ). This approach ensured we captured potential correlations while maintaining the temporal resolution necessary to detect rapid variability.
@@
To further test the robustness of the observed correlations and minimize potential biases from individual data points or small subsets, we performed a jackknife analysis on each component. This analysis, which involves iteratively removing one data point and recalculating the correlation, demonstrated consistent results. The mean correlation coefficients and p-values obtained from the jackknife analysis were in good agreement with those presented in Tables 8 and 9 in all cases. Also, to assess the reliability of the presented correlations, we adopted the methodology described in Alexander (1997). We conducted 10,000 Monte Carlo simulations, generating simulated datasets with fluxes drawn randomly from within their respective error ranges. Spearman rank correlations were then computed for each simulated dataset. Subsequently, we performed Fisherās z-transform on the resulting correlation coefficients, calculating the median and standard deviation in the z-space. The median was then back-transformed to obtain a correlation coefficient, and asymmetric errors were estimated. The last column of Tables 8 and 9 present the correlation coefficients and associated errors determined through this Monte Carlo approach, and they are consistent with the Spearman correlations obtained from the original data.
@@
The results of our analysis are presented in Table 8, they reveal significant correlations between the -ray emission and the 43 GHz flux densities of specific jet components found in the BEAM-ME maps. The core (Q0) shows the strongest correlation, with a coefficient () of 0.52. This suggests a possible link between core variability at 43 GHz and roughly 27 of the observed -ray variability, as indicated by a Spearman rank correlation coefficient squared () of 0.27. Similarly, correlations with components Q3 and Q21 suggest that 26 and 18 of the -ray variability, respectively, could be related to the 43 GHz emission in these components during their detection periods. Figure 6 illustrates these relationships between the simultaneous -ray fluxes and the 43 GHz flux densities of the core and components Q3 and Q21. The data points associated with Q8 and Q21 correspond to two distinct time periods when the quasi-stationary component was observed. This component has been extensively studied in the literature prior to this work (e.g., Pauliny-Toth etĀ al. 1987, Jorstad etĀ al. 2001) and before the Fermi era. Intriguingly, during the first epoch, Q8 exhibits no correlation with the gamma-ray variability, while Q21 shows a clear correlation. As illustrated in Fig. 4, numerous moving features intersect with Q21, whereas none pass through Q8. This observation aligns with models proposed in previous studies that attribute -ray flares to shock-shock interactions between moving and stationary features (e.g., Marscher & Gear 1985; Fromm etĀ al. 2011). These findings suggest that the quasi-stationary component may act as a -ray emission region, with its activity being modulated by the occurrence of such interactions.
@@
Figure 6 (right panels) also depicts the relationships between the -ray fluxes and the 15 GHz flux densities of the core (K0) and components K6 and K8 from the MOJAVE maps. Table 9 summarizes the Spearman rank correlation results. The core, once again, exhibits a strong and significant correlation with -ray emission, with a coefficient of 0.62. This implies that roughly 39 of the -ray variations are likely linked to the core 15 GHz activity. Interestingly, a new jet component, K6, emerged near the end of the study period (early 2020). This component shows a significant anti-correlation with the -ray flux, with a coefficient of , suggesting that nearly 40 of the -ray variability during this specific time might be related to the ejection and 15 GHz emission of K6, even in the absence of a simultaneous flare. As suggested by PatiƱo-Ćlvarez etĀ al. (2018), one possible explanation for the observed behavior is that the jet became less transparent to -rays. This change in transparency might be caused by gamma-ray absorption due to triple pair production. The cross section for both the IC scattering and the triple pair production are dependent on the energy of the electrons in the jet, as when the Lorentz factors of individual electrons increase up to a certain threshold, the triple pair production becomes more prominent than the IC. This process would absorb -rays, causing a flux decrease, while at the same time increasing the synchrotron emission, due to the increase in the Lorentz factor, and the production of more leptons within the jet. However, the proximity of K6 to the core as seen in Fig. 2, near the spatial resolution limit, raises the possibility of flux contamination. Additionally, a quasi-stationary component, K8, was detected at approximately 53 pc from the core. Although correlation analysis was feasible with 11 epochs, more observations are essential to determine this relationship.
@@
This study offers novel insights into the -ray emission in blazars. Our findings reveal the presence of multiple distinct regions within the jet, with each contributing to the high-energy emission. Notably, some of these regions exhibit non-stationary behavior and display temporal variability. Furthermore, we successfully located these emission regions within the jet structure. Previous multi-frequency studies of 3C 454.3 have revealed strong correlations between -ray and optical emissions (Ogle etĀ al. 2011; León-Tavares etĀ al. 2013; VillicaƱa-Pedraza etĀ al. 2016; Nalewajko etĀ al. 2019; Amaya-AlmazĆ”n etĀ al. 2021; Fan etĀ al. 2021). These correlations strongly suggest that rays in blazars originate from leptonic processes rather than hadronic ones, as no known hadronic mechanism can simultaneously generate rays and significantly impact optical synchrotron emission. To distinguish between SSC and EIC as the dominant leptonic process, determining the location of the emitting region within the jet is crucial. For regions located several parsecs away from the central engine, the density of external seed photons is typically lower than the synchrotron photon density, thus favoring SSC as the primary emission mechanism. This is supported by findings in 3C 279, where PatiƱo-Ćlvarez etĀ al. (2019) concluded that approximately 33% of the observed -ray emission originates from SSC processes within a particular jet component during its the observable period. In the core region of the blazar, while we can confirm a leptonic origin for the gamma-ray emission, differentiating between SSC and EIC remains challenging. However, it is generally accepted that sufficiently energetic electrons exist in the core to power the observed -ray emission. For non-core components (i.e., those located farther from the central parsec), the correlation observed in simultaneous data suggests that the -ray emission mechanism is SSC. The electron energy distribution within a knot is such that while lower-energy electrons dominate the radio band, higher-energy electrons capable of producing gamma-rays can co-exist. These findings can guide the development of more comprehensive theoretical models that might account for the observed variability and spatial distribution of high-energy emissions.
@@ 4 Summary and conclusions
In this study, we have analyzed the jet behavior and structure of the blazar 3C 454.3 using radio data from the VLBA-BU-BLAZAR, BEAM-ME, and MOJAVE monitoring programs (at frequencies of 43 GHz and 15 GHz). We investigated potential -ray emission regions within the jet components identified in the VLBA maps. This included measuring the distances between these regions and the radio core and exploring possible emission mechanisms. Additionally, we examined whether these regions account for all the -ray emission or, as in the case of 3C 279 (PatiƱo-Ćlvarez etĀ al., 2019), multiple regions are needed to describe it. Our key findings include the following:
@@
Both 43 GHz and 15 GHz core emissions strongly correlate with -ray emission. We observed a robust correlation between the 43 GHz core emission and -ray emission, with a correlation coefficient of 0.52. This indicates that approximately 27 of the variability in the -ray emission is related to the variability in the core radio emission. Similarly, the flux of the 15 GHz core exhibits a strong and statistically significant correlation with the -ray emission, contributing approximately 40 to the high-energy emission.
@@
The quasi-stationary component Q21 in the 43 GHz maps, visible between 2016 and 2021, at a projected distance of 4.6 pc (0.6 mas) strongly correlates with the -ray flux, implying that it may contribute up to 18% of the high-energy emission. However, the quasi-stationary component observed during a different time interval, Q8, does not correlate. This variability in the correlation suggests that quasi-stationary components may exhibit intermittent gamma-ray emission.
@@
At 43 GHz, we detected a mobile component whose flux exhibits a robust correlation with the -ray flux, contributing 26 of the high-energy emission between 2010.18 and 2011.16. This component, denoted as Q3, demonstrates an apparent velocity of c, observed at a projected distance ranging from 0.8 to 2.3 pc (0.1 to 0.3 mas).
@@
For -ray emission regions located farther than one parsec from the central black hole (e.g., components Q21 and Q3), our findings strongly suggest SSC as the primary mechanism responsible for the high-energy radiation. This conclusion is supported by the observed correlation between the radio variability in these regions and the -ray emission, indicating a nearly simultaneous variability.
@@
In the 15 GHz maps, a new jet component, K6, appeared in early 2020. It strongly anti-correlates with -ray flux (correlation coefficient -0.65), suggesting a potential link to around 40 of the -ray variability. However, the proximity of K6 to the core might distort its measured flux. Another component, K8, shows a similar anti-correlation with -rays, but we believe more data is needed to confirm this behavior.
@@
The observed anti-correlations in the two 15 GHz components may be attributed to an increased Lorentz factor in the electrons of the -ray emission region, which can enhance electron-positron triple pair production and lead to greater -ray absorption and a subsequent decrease in observed gamma-ray flux, as suggested by PatiƱo-Ćlvarez etĀ al. (2018). This would also increase the synchrotron emission because the energy and number of the leptons are increasing, thus creating the observed anti-correlation.
Our findings are valuable on two fronts. First, they provide crucial data for refining theoretical models of blazar jets. Second, they open new avenues for investigation by pinpointing specific regions within the jet for further study. This comprehensive understanding might lead to a deeper grasp of the complex processes at play in blazars.